The rawfonts package

Alan Jeffrey

v0.01

Overview

The rawfonts L^{A}TEX 2ε package is used to provide emulation of L^{A}TEX 2.09 documents which used low-level commands such as \tenrm. If you say:

    \usepackage{rawfonts}

then over sixty fonts such as \tenrm will be loaded into L^{A}TEX. This has a large overhead, for example this document uses:

    8221 words of font info for 30 fonts

but when run with the rawfonts package it uses:

    19294 words of font info for 73 fonts

That is rawfonts can cause your document to load twice as many fonts. (This overhead is why L^{A}TEX 2ε does not define \tenrm and friends by default.)

If you want to only load a small number of fonts, you can use the only option, for example to only load \tenrm and \tensf:

    \usepackage[only,tenrm,tensf]{rawfonts}

The rawfonts package is intended mainly for use with L^{A}TEX 2.09 documents, and you might want to have the package loaded every time you use L^{A}TEX 2ε in compatibility mode. In this case, you should add the line:

    \RequirePackage{rawfonts}

to your latex209.cfg L^{A}TEX 2.09 compatibility configuration file.

Implementation

The driver for the documentation you’re now reading.

1 ⟨∗driver⟩

 2 \documentclass{ltxdoc}
 3 \begin{document}
 4 \DocInput{rawfonts.dtx}
 5 \end{document}

6 ⟨/driver⟩

This is a L^{A}TEX 2ε package.

7 ⟨∗package⟩

 8 \NeedsTeXFormat{LaTeX2e}
 9 \ProvidesPackage{rawfonts}
10 [1994/05/08 Low-level LaTeX 2.09 font compatibility]

The rawfonts package uses the somedefs package.

11 \RequirePackage{somedefs}

By default, all the fonts are loaded, but the only option says that only the ones specified as options should be.

12 \UseAllDefinitions
13 \DeclareOption{only}{\UseSomeDefinitions}
14 \DeclareOption*{\UseDefinition{\CurrentOption}}
15 \ProcessOptions

The rest of the code loads fonts. Five point:

16 \ProvidesDefinition{\DeclareFixedFont{\fivrm}{OT1}{cmr}{m}{n}{\@vpt}}
17 \ProvidesDefinition{\DeclareFixedFont{\fivmi}{OML}{cmm}{m}{it}{\@vpt}}
18 \ProvidesDefinition{\DeclareFixedFont{\fivsy}{OMS}{cmsy}{m}{n}{\@vpt}}
19 \ProvidesDefinition{\DeclareFixedFont{\fivly}{U}{lasy}{m}{n}{\@vpt}}

Six point:

20 \ProvidesDefinition{\DeclareFixedFont{\sixrm}{OT1}{cmr}{m}{n}{\@vipt}}
21 \ProvidesDefinition{\DeclareFixedFont{\sixmi}{OML}{cmm}{m}{it}{\@vipt}}
22 \ProvidesDefinition{\DeclareFixedFont{\sixsy}{OMS}{cmsy}{m}{n}{\@vipt}}
23 \ProvidesDefinition{\DeclareFixedFont{\sixly}{U}{lasy}{m}{n}{\@vipt}}

Seven point:

24 \ProvidesDefinition{\DeclareFixedFont{\sevrm}{OT1}{cmr}{m}{n}{\@viipt}}
25 \ProvidesDefinition{\DeclareFixedFont{\sevmi}{OML}{cmm}{m}{it}{\@viipt}}
26 \ProvidesDefinition{\DeclareFixedFont{\sevsy}{OMS}{cmsy}{m}{n}{\@viipt}}
27 \ProvidesDefinition{\DeclareFixedFont{\sevit}{OT1}{cmr}{m}{it}{\@viipt}}
28 \ProvidesDefinition{\DeclareFixedFont{\sevly}{U}{lasy}{m}{n}{\@viipt}}

Eight point:

29 \ProvidesDefinition{\DeclareFixedFont{\egtrm}{OT1}{cmr}{m}{n}{\@viiipt}}
30 \ProvidesDefinition{%
31  \DeclareFixedFont{\egtmi}{OML}{cmm}{m}{it}{\@viiipt}}
32 \ProvidesDefinition{%
33  \DeclareFixedFont{\egtsy}{OMS}{cmsy}{m}{n}{\@viiipt}}
34 \ProvidesDefinition{%

35  \DeclareFixedFont{\egtit}{OT1}{cmr}{m}{it}{\@viiipt}}
36 \ProvidesDefinition{\DeclareFixedFont{\egtly}{U}{lasy}{m}{n}{\@viiipt}}

Nine point:

37 \ProvidesDefinition{\DeclareFixedFont{\ninrm}{OT1}{cmr}{m}{n}{\@ixpt}}
38 \ProvidesDefinition{\DeclareFixedFont{\ninmi}{OML}{cmm}{m}{it}{\@ixpt}}
39 \ProvidesDefinition{\DeclareFixedFont{\ninsy}{OMS}{cmsy}{m}{n}{\@ixpt}}
40 \ProvidesDefinition{\DeclareFixedFont{\ninit}{OT1}{cmr}{m}{it}{\@ixpt}}
41 \ProvidesDefinition{\DeclareFixedFont{\ninbf}{OT1}{cmr}{bx}{n}{\@ixpt}}
42 \ProvidesDefinition{\DeclareFixedFont{\nintt}{OT1}{cmtt}{m}{n}{\@ixpt}}
43 \ProvidesDefinition{\DeclareFixedFont{\ninly}{U}{lasy}{m}{n}{\@ixpt}}

Ten point:

44 \ProvidesDefinition{\DeclareFixedFont{\tenrm}{OT1}{cmr}{m}{n}{\@xpt}}
45 \ProvidesDefinition{\DeclareFixedFont{\tenmi}{OML}{cmm}{m}{it}{\@xpt}}
46 \ProvidesDefinition{\DeclareFixedFont{\tensy}{OMS}{cmsy}{m}{n}{\@xpt}}
47 \ProvidesDefinition{\DeclareFixedFont{\tenit}{OT1}{cmr}{m}{it}{\@xpt}}
48 \ProvidesDefinition{\DeclareFixedFont{\tensl}{OT1}{cmr}{m}{sl}{\@xpt}}
49 \ProvidesDefinition{\DeclareFixedFont{\tenbf}{OT1}{cmr}{bx}{n}{\@xpt}}
50 \ProvidesDefinition{\DeclareFixedFont{\tentt}{OT1}{cmtt}{m}{n}{\@xpt}}
51 \ProvidesDefinition{\DeclareFixedFont{\tensf}{OT1}{cmss}{m}{n}{\@xpt}}
52 \ProvidesDefinition{\DeclareFixedFont{\tenly}{U}{lasy}{m}{n}{\@xpt}}
53 \ProvidesDefinition{\DeclareFixedFont{\tenex}{OMX}{cmex}{m}{n}{\@xpt}}

Eleven point:

54 \ProvidesDefinition{\DeclareFixedFont{\elvrm}{OT1}{cmr}{m}{n}{\@xipt}}
55 \ProvidesDefinition{\DeclareFixedFont{\elvmi}{OML}{cmm}{m}{it}{\@xipt}}
56 \ProvidesDefinition{\DeclareFixedFont{\elvsy}{OMS}{cmsy}{m}{n}{\@xipt}}
57 \ProvidesDefinition{\DeclareFixedFont{\elvit}{OT1}{cmr}{m}{it}{\@xipt}}
58 \ProvidesDefinition{\DeclareFixedFont{\elvsl}{OT1}{cmr}{m}{sl}{\@xipt}}
59 \ProvidesDefinition{\DeclareFixedFont{\elvbf}{OT1}{cmr}{bx}{n}{\@xipt}}
60 \ProvidesDefinition{\DeclareFixedFont{\elvtt}{OT1}{cmtt}{m}{n}{\@xipt}}
61 \ProvidesDefinition{\DeclareFixedFont{\elvsf}{OT1}{cmss}{m}{n}{\@xipt}}
62 \ProvidesDefinition{\DeclareFixedFont{\elvly}{U}{lasy}{m}{n}{\@xipt}}

Twelve point:

63 \ProvidesDefinition{\DeclareFixedFont{\twlrm}{OT1}{cmr}{m}{n}{\@xiipt}}
64 \ProvidesDefinition{\DeclareFixedFont{\twlmi}{OML}{cmm}{m}{it}{\@xiipt}}
65 \ProvidesDefinition{\DeclareFixedFont{\twlsy}{OMS}{cmsy}{m}{n}{\@xiipt}}
66 \ProvidesDefinition{\DeclareFixedFont{\twlit}{OT1}{cmr}{m}{it}{\@xiipt}}
67 \ProvidesDefinition{\DeclareFixedFont{\twlsl}{OT1}{cmr}{m}{sl}{\@xiipt}}
68 \ProvidesDefinition{\DeclareFixedFont{\twlbf}{OT1}{cmr}{bx}{n}{\@xiipt}}
69 \ProvidesDefinition{\DeclareFixedFont{\twltt}{OT1}{cmtt}{m}{n}{\@xiipt}}
70 \ProvidesDefinition{\DeclareFixedFont{\twlsf}{OT1}{cmss}{m}{n}{\@xiipt}}
71 \ProvidesDefinition{\DeclareFixedFont{\twlly}{U}{lasy}{m}{n}{\@xiipt}}

Fourteen point:

72 \ProvidesDefinition{\DeclareFixedFont{\frtnrm}{OT1}{cmr}{m}{n}{\@xivpt}}
73 \ProvidesDefinition{%
74  \DeclareFixedFont{\frtnmi}{OML}{cmm}{m}{it}{\@xivpt}}
75 \ProvidesDefinition{%
76  \DeclareFixedFont{\frtnsy}{OMS}{cmsy}{m}{n}{\@xivpt}}

77 \ProvidesDefinition{%
78  \DeclareFixedFont{\frtnbf}{OT1}{cmr}{bx}{n}{\@xivpt}}
79 \ProvidesDefinition{\DeclareFixedFont{\frtnly}{U}{lasy}{m}{n}{\@xivpt}}

Seventeen point:

80 \ProvidesDefinition{%
81  \DeclareFixedFont{\svtnrm}{OT1}{cmr}{m}{n}{\@xviipt}}
82 \ProvidesDefinition{%
83  \DeclareFixedFont{\svtnmi}{OML}{cmm}{m}{it}{\@xviipt}}
84 \ProvidesDefinition{%
85  \DeclareFixedFont{\svtnsy}{OMS}{cmsy}{m}{n}{\@xviipt}}
86 \ProvidesDefinition{%
87  \DeclareFixedFont{\svtnbf}{OT1}{cmr}{bx}{n}{\@xviipt}}
88 \ProvidesDefinition{\DeclareFixedFont{\svtnly}{U}{lasy}{m}{n}{\@xviipt}}

Twenty point:

89 \ProvidesDefinition{\DeclareFixedFont{\twtyrm}{OT1}{cmr}{m}{n}{\@xxpt}}
90 \ProvidesDefinition{\DeclareFixedFont{\twtymi}{OML}{cmm}{m}{it}{\@xxpt}}
91 \ProvidesDefinition{\DeclareFixedFont{\twtysy}{OMS}{cmsy}{m}{n}{\@xxpt}}
92 \ProvidesDefinition{\DeclareFixedFont{\twtyly}{U}{lasy}{m}{n}{\@xxpt}}

Twenty-five point:

93 \ProvidesDefinition{\DeclareFixedFont{\twfvrm}{OT1}{cmr}{m}{n}{\@xxvpt}}

That’s it.

94 ⟨/package⟩